# Compositional Difference-in-Differences for categorical outcomes

### v1

@@ Abstract.
In difference-in-differences (DiD) settings with categorical outcomes, treatment effects often operate on both total quantities (e.g., voter turnout) and category shares (e.g., vote distribution across parties). In this context, linear DiD models can be problematic: they suffer from scale dependence, may produce negative counterfactual quantities, and are inconsistent with discrete choice theory. We propose compositional DiD (CoDiD), a new method that identifies counterfactual categorical quantities, and thus total levels and shares, under a parallel growths assumption. The assumption states that, absent treatment, each category’s size grows or shrinks at the same proportional rate in treated and control groups. In a random utility framework, we show that this implies parallel evolution of relative preferences between any pair of categories. Analytically, we show that it also means the shares are reallocated in the same way in both groups in the absence of treatment. Finally, geometrically, it corresponds to parallel trajectories (or movements) of probability mass functions of the two groups in the probability simplex under Aitchison geometry. We extend CoDiD to (i) derive bounds under relaxed assumptions, (ii) handle staggered adoption, and (iii) propose a synthetic DiD analog. We illustrate the method’s empirical relevance through two applications: first, we examine how early voting reforms affect voter choice in U.S. presidential elections; second, we analyze how the regional greenhouse gas initiative (RGGI) affected the composition of electricity generation across sources such as coal, natural gas, nuclear, and renewables.
Keywords: categorical outcomes, difference-in-differences, causal inference, identification, changes-in-changes, compositional data.
@@
Corresponding address: oib5044@psu.edu, Department of Economics, The Pennsylvania State University, University Park, PA 16802. We are grateful to Marc Henry, Andres Aradillas-Lopez, Sung Jae Jun, and Keisuke Hirano for their guidance on this project. We are also grateful to the participants of the Penn State Econometrics Seminar, the Penn State Student Econometrics Workshop, and the RESA Alumni Conference 2025 for their valuable comments and suggestions. All errors are mine.
@@ Introduction
The classical difference-in-differences (DiD) framework identifies the average treatment effect on the treated under the parallel trends assumption for scalar outcomes. However, in many applications, the outcome of interest is categorical, and the notion of an average level is generally undefined. Examples include voting decision (democrat, republican, others), occupational sector (agriculture, manufacturing, and services) or field of study (stem, humanities, business, or others). For categorical outcomes, the key quantitative objects are the category-specific quantities, which jointly determine both the overall total quantity and the corresponding probability mass function across categories. In this case, we would like to analyze both absolute changes (how many units are added or removed in absolute terms for each category, and overall) and relative shifts (how the shares are reallocated across categories) in a consistent way. For example, a voting policy can differently affect both the total number of voters (turnout) and the share of votes for each party. A minimum wage policy can change both the number of people across all employment sectors and the probability of being in each of them. An education reform may reallocate students between fields, but policymakers may also care about the total number of new graduates overall.
@@
The classical linear parallel trends assumption, which imposes a linear evolution in outcome levels, may be inappropriate in this context for several reasons already noted in the literature (see Puhani (2012); Wooldridge (2023)). A naive strategy would be to apply linear parallel trends separately to the raw quantities of each category, obtain counterfactual quantities, and then normalize them to form a counterfactual distribution. However, this approach is problematic for three main reasons. First, applying linear parallel trends to raw quantities is scale-dependent, as it ignores differences in group size. Transferring the same additive change from a large control group to a smaller treated group can distort counterfactual quantities, exaggerating or understating effects, and thus producing misleading totals and distributions. Second, this method can yield negative counterfactual quantities when a negative shift in the control group outweighs the pre-treatment value in the treated group. Finally, it can imply choice dynamics that lack any behavioral or structural justification from economic theory regarding how choices across categories evolve.
@@
On top of that, applying parallel trends directly to probabilities is theoretically possible but often yields invalid counterfactuals, with values below zero or exceeding one. The resulting counterfactual distribution can fall outside the probability simplex (the set of all vectors of positive numbers summing to one), rendering them invalid (see figure 1). Most existing nonlinear methods for categorical distributions require a natural ordering of the support because they rely on cumulative distribution functions, which are undefined for unordered outcomes. For such outcomes, only the probability mass function is well-defined, and importantly, these methods typically do not provide counterfactuals for the raw quantities.
@@
Ideally, we would like to consider a nonlinear parallel trends assumption on quantities that has a clear economic meaning, resolves the scale-dependence problem, and ensures that the evolution of shares remains consistent with standard discrete choice theory. This paper introduces Compositional difference-in-differences (CoDiD), a new framework for analyzing categorical outcomes. The method is built on a nonlinear parallel trends assumption, which we call parallel growths, which allows us to point-identify the counterfactual categorical quantity, and thus the counterfactual total and probability mass functions within a DiD framework. We do so while addressing all the challenges outlined above. In addition, we define two treatment effect measures tailored to this context. The growth treatment effect on the treated (GTT) quantifies how each category’s absolute size changes as a result of treatment, and also how the total quantity has evolved as a result of treatment. The second is the compositional treatment effect on the treated (CTT), which captures how the probability mass is reallocated across categories in response to treatment.
@@
We begin with the canonical $2\times 2$ DiD setting. The parallel growths assumption is a parallel trends of log-transformed categorical quantities, meaning that the absolute size of each category would grow or shrink at the same proportional rate in both treated and control groups. We first show that this directly implies parallel trends in the log-odds of category shares. Economically, this connects to the random utility model (McFadden (1972)), where log-odds represent differences in utilities and reflect relative preferences. Under this interpretation, parallel growths mean that, absent treatment, the evolution of relative preferences between any pair of categories would be similar in both groups. Therefore, the way people chose between categories would not change if there is no treatment. Also, without relying on a random utility framework, we can derive another economic meaning. In the control group, category shares naturally evolve over time, with some categories gaining mass at the expense of others. The assumption implies that, in the absence of treatment, the treatment group would have experienced the same reallocation pattern observed in the control group. This interpretation allows us to view the causal effect as the extent to which the treatment affects the way probability mass is redistributed across categories. Finally, in geometric terms, each probability mass function (PMF) corresponds to a point in the simplex, and its evolution traces a path within this space. The parallel growths assumption means that, in the absence of treatment, the treated group’s counterfactual path would be parallel to the control group’s path under Aitchison geometry, the standard geometry for the simplex in compositional data analysis (Aitchison (1982); Egozcue et al. (2003)).
@@
We also discuss the connections and key differences between our approach and the multinomial logistic two-way fixed effects model. In the simple $2\times 2$ case, we show that, when focusing solely on category shares, both methods yield the same counterfactual shares. The main distinction lies in the fact that our approach also recovers counterfactual quantities, which cannot be obtained directly from the logistic model. Moreover, our framework introduces a new treatment effect parameter that offers a more natural and interpretable measure than the change in log-odds implied by the logistic specification. Even when the analysis is limited to shares, the differences become more pronounced when extending the framework to multiple time periods (e.g. staggered adoption case) and relaxations of the identifying asusmption. Our approach provides a more flexible and easily extendable foundation for such generalizations. Another comparison is made with the DiD and the changes-in-changes (CiC) frameworks. While DiD focuses on expectations (points on the real line) and CiC on cumulative distribution functions (points in a space of CDFs), our method focuses on PMFs (points in the probability simplex). The underlying identification logic is consistent across all three methods: each characterizes the control group’s evolution via a translation mapping and applies that same mapping to the treated group’s initial state to recover its counterfactual. The fundamental distinction, therefore, lies not in the core logic but in the object being transported, expectations, CDFs, or PMFs, and the geometry that defines the translation. From this unified perspective, the CoDiD assumption emerges naturally as the DiD/CiC analogue for probability mass functions.
@@
In a more general framework with multiple time periods, the parallel growths assumption can be assessed visually by plotting the log-transformed quantities. If the trends appear non-parallel in pre-treatment periods, the assumption becomes less credible. In this case, we first propose a set of relaxations of the assumption in the spirit of Ban and Kédagni (2022). Specifically, we bound the counterfactual log differences between groups by restricting their change to lie within the convex hull of observed changes across pre-treatment periods. This approach produces bounds for both counterfactual quantities and counterfactual distributions. When multiple potential control groups are available, we develop a synthetic DiD (see Arkhangelsky et al. (2021)) analog, synthetic CoDiD, that merges the flexibility of synthetic control methods with the causal logic of difference-in-differences. This extension constructs a weighted combination of control units whose pre-treatment log-trajectory closely matches that of the treated unit, while preserving the parallel growths structure in the post-treatment period. The resulting control group not only improves pre-treatment fit but also maintains interpretability within the CoDiD framework. Finally, we generalize the methodology to settings with staggered treatment adoption, where units receive treatment at different points in time. We follow the same assumptions as in Callaway and Sant’Anna (2021), propose two types of parallel growths assumption in this case, and derive the corresponding counterfactual in each case.
@@
As applications, we first examine the impact of early voting programs on the distribution of voter support across political parties in U.S. presidential elections. These programs, allowing voters to cast ballots before election day, either in person or by mail, have been the subject of intense partisan debate. We revisit this question by analyzing both total turnout and the compositional shift in voter preferences. As expected, the treatment increased overall turnout and the number of votes for all parties. However, our compositional treatment effect on the Treated (CTT) reveals a nuanced reallocation: a previously indifferent voter (assigning equal initial weight to all parties) would, after the treatment, shift support more toward alternative parties, with both major parties losing relative appeal, Republicans more so than Democrats. Secondly, we analyze the Regional Greenhouse Gas Initiative (RGGI) and its effects on the mix of electricity generation sources in participating states. RGGI, a cap-and-trade program designed to reduce carbon emissions from the power sector, is expected to shift the composition of electricity generation across coal, natural gas, nuclear, and renewable sources by promoting cleaner technologies and discouraging carbon-intensive ones. We assess how RGGI affected the size of each source in the electricity generation process over time. We find that RGGI sharply cut fossil fuel use, especially coal, by shifting generation to gas and nuclear rather than boosting renewables, underscoring the need for complementary policies to accelerate a truly renewable-driven energy transition.
@@
First, it contributes to the extensive DiD literature. Classical DiD focuses on scalar outcomes, estimating treatment effects on the mean, with surveys and applications summarized in Lechner (2011); Cafri et al. (2019); De Chaisemartin and d’Haultfoeuille (2023); Roth et al. (2023); Baker et al. (2025). Recent works have extended DiD to more complex settings: Callaway and Sant’Anna (2020) develops methods for staggered treatment adoption, Arkhangelsky et al. (2021) combines DiD with synthetic controls, and Manski and Pepper (2018); Rambachan and Roth (2020); Ban and Kédagni (2022) relax the parallel trends assumption. A second wave of research extends DiD to entire outcome distributions. Early contributions include quantile DiD Meyer et al. (1990) and the changes-in-changes (CiC) approach Athey and Imbens (2006), which identifies counterfactual distributions non-parametrically. Subsequent work has generalized these ideas to multivariate outcomes Torous et al. (2021), settings where identifying assumptions apply to cumulative distribution functions Havnes and Mogstad (2015); Roth and Sant’Anna (2021), copula-based approaches Callaway et al. (2018); Callaway and Li (2019); Ghanem et al. (2023), and methods using characteristic functions Bonhomme and Sauder (2011). However, these methods do not directly address categorical outcome distributions, which are probability mass functions. Graves et al. (2022) proposes a DiD for categorical outcomes while relying on parallel trends for proportions, which we explained earlier is not appropriate in general. Zhou et al. (2025) extend the difference-in-differences framework to non-Euclidean data, including probability mass functions, by employing Fréchet means and geodesic transport. While their approach is mathematically elegant, the identifying assumptions are primarily geometric and offer limited economic interpretability. In contrast, our framework delivers both counterfactual quantities and probability mass functions that retain economic intuition while incorporating a geometric perspective specifically designed for categorical outcomes. Another related paper is Tchetgen Tchetgen et al. (2024), who propose a general DiD framework applicable to count data in the binomial setting. Our analysis differs by focusing on the multinomial case, allowing for richer categorical structures and more general forms of distributional change. Second, the paper contributes to the compositional data analysis (CoDA) literature. CoDA, pioneered by Aitchison (1982, 1990, 1992, 2002), provides tools for analyzing data constrained to the simplex. Subsequent developments Egozcue et al. (2003); Billheimer et al. (2001); Barceló-Vidal et al. (2001) refine transformations and models that respect the geometric structure of compositional data. More recently, Arnold et al. (2020) connected CoDA to causal inference. We also combine compositional data methods with econometric identification strategies to analyze causal effects when outcomes are probability mass functions, providing a bridge between DiD and compositional statistics.

### v2

@@ Abstract.
In difference-in-differences (DiD) settings with categorical outcomes, such as voting, occupation, or major choices, treatments often affect both total counts (e.g., turnout) and category shares (e.g., vote shares). Traditional linear DiD models can yield invalid counterfactuals in this context (e.g. negative values ) and lack compatibility with standard discrete choice models. I propose Compositional DiD (CoDiD), which identifies counterfactual totals and shares under a parallel growths assumption: absent treatment, each category’s size grows or shrinks at the same proportional rate in treated and control groups. I show that under a random utility model, this is equivalent to parallel trends in expected utilities, i.e., the change in average latent desirability for each alternative is identical across groups. Consequently, relative preferences (e.g., how individuals prefer Democrat vs. Republican) evolve in parallel, and counterfactual distributions follow parallel trajectories in the probability simplex. I extend CoDiD to (i) derive bounds under relaxed assumptions, (ii) accommodate staggered treatment timing, and (iii) construct a synthetic DiD analog. I illustrate the method’s empirical relevance through two applications: first, I examine how early voting reforms affect voter choice in U.S. presidential elections; second, I analyze how the Regional Greenhouse Gas Initiative (RGGI) affected the electricity mix in participating states.
Keywords: categorical outcomes, difference-in-differences, causal inference, compositional data, voting, rggi.
@@
Corresponding address: oib5044@psu.edu, Department of Economics, The Pennsylvania State University, University Park, PA 16802. I am grateful to Marc Henry, Andres Aradillas-Lopez, Sung Jae Jun, and Keisuke Hirano for their guidance on this project. I am also grateful to the participants of the Penn State Econometrics Seminar, the Penn State Student Econometrics Workshop, and the RESA Alumni Conference 2025 for their valuable comments and suggestions. All errors are mine.
@@ Introduction
The classical difference-in-differences (DiD) framework identifies the average treatment effect on the treated under the parallel trends assumption for scalar outcomes. However, in many applications, the outcome of interest is categorical, and the notion of an average level is generally undefined. Examples include voting decision (democrat, republican, others), occupational sector (agriculture, manufacturing, and services) or field of study (stem, humanities, business, or others). For categorical outcomes, the key quantitative objects are the category-specific quantities, which jointly determine both the overall total quantity and the corresponding probability mass function across categories. In this case, one would like to analyze both absolute changes (how many units are added or removed in absolute terms for each category, and overall) and relative changes (how the shares are reallocated across categories) in a consistent way. For example, a voting policy can differently affect both the total number of voters (turnout) and the reallocation of vote shares for each party. A minimum wage policy can change both the number of people across all employment sectors and the probability of being in each of them. An education reform may reallocate students between fields, but policymakers may also care about the total number of new graduates overall.
@@
The classical linear parallel trends assumption, which imposes a linear evolution in outcome levels, may be inappropriate in this context for several reasons already noted in the literature (see Puhani (2012); Wooldridge (2023)). A naive strategy would be to apply linear parallel trends separately to the raw quantities of each category, obtain counterfactual quantities, and then normalize them to form a counterfactual distribution. However, this approach is problematic for three main reasons. First, applying linear parallel trends to raw quantities is scale-dependent, as it ignores differences in group size. Transferring the same additive change from a large control group to a smaller treated group can distort counterfactual quantities, exaggerating or understating effects, and thus producing misleading totals and distributions. Second, this method can yield negative counterfactual quantities when a negative shift in the control group outweighs the pre-treatment value in the treated group. Finally, it would imply choice dynamics that lack any behavioral or structural justification from economic theory regarding how choices across categories evolve.
@@
On top of that, applying parallel trends directly to probabilities is theoretically possible but often yields invalid counterfactuals, with values below zero or exceeding one. The resulting counterfactual distribution can fall outside the probability simplex (the set of all vectors of positive numbers summing to one), rendering them invalid (see figure 1). Most existing nonlinear methods for categorical distributions require a natural ordering of the support because they rely on cumulative distribution functions, which are undefined for unordered outcomes. For such outcomes, only the probability mass function is well-defined, and importantly, these methods typically do not directly provide counterfactuals for the total quantities.
@@
Ideally, one would like to consider a nonlinear parallel trends assumption on categorical quantities that has a clear natural meaning, resolves the scale-dependence problem, and ensures that the evolution of shares remains consistent with standard discrete choice theory. This paper introduces Compositional difference-in-differences (CoDiD), a new framework for analyzing categorical outcomes in a DiD setting. The method is built on a nonlinear parallel trends assumption, which I call parallel growths, which allows us to point-identify the counterfactual categorical quantity, and thus the counterfactual total and probability mass functions within a DiD framework. I do so while addressing all the challenges outlined above. In addition, I define two treatment effect measures tailored to this context. The growth treatment effect on the treated (GTT) quantifies how each category’s absolute size changes as a result of treatment, and also how the total quantity has evolved as a result of treatment. The second is the compositional treatment effect on the treated (CTT), which captures how the probability mass is reallocated across categories in response to treatment.
@@
I begin with the canonical $2\times 2$ DiD setting. The parallel growths assumption is a parallel trends of log-transformed categorical quantities, meaning that the absolute size of each category would grow or shrink at the same proportional rate in both treated and control groups. Using a standard random utility model, I show that this assumption is equivalent to assuming parallel trends of latent utilities for each category. This means that in the absence of treatment, both groups would have experienced the same evolution in how desirable each option is on average. For example, if in the control group, the expected utility for voting democrat increase, the same movement will happen in treated group in absence of treatment. The implied evolution of the shares is a parallel trends in the log-odds of category shares. In the random utility model, log-odds capture utiliy differences, therefore relative preferences. Under this interpretation, parallel growths mean that, absent treatment, the evolution of relative preferences between any pair of categories would be similar in both groups. In other words, the evolution of the way people chose between categories in the control group will be similar to what happens in the treated group. Under this interpretation, we can view the causal effect as the way the treatment shift the way categories are desired.
Furthermore, without relying on a random utility framework, I derive another economic implication of parallel growths. In the control group, category shares naturally evolve over time, with some categories gaining mass at the expense of others. The assumption implies that, in the absence of treatment, the treatment group would have experienced the same reallocation pattern observed in the control group. This interpretation allows us to view the causal effect as the extent to which the treatment change the way probability mass is redistributed across categories. Finally, in geometric terms, each probability mass function (PMF) corresponds to a point in the simplex, and its evolution traces a path within this space. The parallel growths assumption means that, in the absence of treatment, the treated group’s counterfactual path would be parallel to the control group’s path under Aitchison geometry, the standard geometry for the simplex in compositional data analysis (Aitchison (1982); Egozcue et al. (2003)).
@@
I also discuss the connections and key differences between my approach and the multinomial logistic two-way fixed effects model. In the simple $2\times 2$ case, I show that, when focusing solely on category shares, both methods yield the same counterfactual shares. The main distinction lies in the fact that my approach also recovers counterfactual quantities, which cannot be obtained directly from the logistic model. Moreover, my framework introduces a new treatment effect parameter that offers a more natural and interpretable measure than the change in log-odds implied by the logistic specification. Even when the analysis is limited to shares, the differences become more pronounced when extending the framework to multiple time periods (e.g. staggered adoption case) and relaxations of the identifying assumption. My approach provides a more flexible and easily extendable foundation for such generalizations. Another comparison is made with the DiD and the changes-in-changes (CiC) frameworks. While DiD targets on expectations (points on the real line) and CiC, cumulative distribution functions (points in a space of CDFs), my method also targets probability mass functions (points in the probability simplex). The underlying identification logic is consistent across all three methods: each characterizes the control group’s evolution via a translation mapping and applies that same mapping to the treated group’s initial state to recover its counterfactual. The fundamental distinction, therefore, lies not in the core logic but in the object being transported, expectations, CDFs, or PMFs, and the geometry that defines the translation. From this unified perspective, the CoDiD emerges naturally as the DiD/CiC analogue for probability mass functions.
@@
In a more general framework with multiple time periods, the parallel growths assumption can be assessed visually by plotting the log-transformed quantities. If the trends appear non-parallel in pre-treatment periods, the assumption becomes less credible. In this case, I first propose a set of relaxations of the assumption in the spirit of Ban and Kédagni (2022). Specifically, I bound the counterfactual log differences between groups by restricting their change to lie within the convex hull of observed changes across pre-treatment periods. This approach produces bounds for both counterfactual quantities and counterfactual distributions. When multiple potential control groups are available, I develop a synthetic DiD (see Arkhangelsky et al. (2021)) analog, synthetic CoDiD, that merges the flexibility of synthetic control methods with the causal logic of difference-in-differences. This extension constructs a weighted combination of control units whose pre-treatment log-trajectory closely matches that of the treated unit, while preserving the parallel growths structure in the post-treatment period. The resulting control group not only improves pre-treatment fit but also maintains interpretability within the CoDiD framework. Finally, I generalize the methodology to settings with staggered treatment adoption, where units receive treatment at different points in time. I follow the same assumptions as in Callaway and Sant’Anna (2021), propose two types of parallel growths assumption in this case, and derive the corresponding counterfactual in each case.
@@
As applications, I first examine the impact of early voting programs on the distribution of voter support across political parties in U.S. presidential elections. These programs, allowing voters to cast ballots before election day, either in person or by mail, have been the subject of intense partisan debate. I revisit this question by analyzing both total turnout and the compositional shift in voter preferences. I consider Maryland and New Jersey, which introduced early voting before 2008 in treated group, while Pennsylvania and New York serve as the control group. As expected, the treatment increased overall turnout ( increase of $4.4\%$) and the number of votes for all parties. However, the compositional treatment effect on the Treated (CTT) reveals a nuanced reallocation: a previously indifferent voter (assigning equal initial weight to all parties) would, after the treatment, shift support more toward alternative parties, with both major parties losing relative appeal, Republicans more so than Democrats. Secondly, I analyze the Regional Greenhouse Gas Initiative (RGGI) and its effects on the mix of electricity generation sources in participating states. RGGI, a cap-and-trade program designed to reduce carbon emissions from the power sector, is expected to shift the composition of electricity generation across coal, natural gas, nuclear, and renewable sources by promoting cleaner technologies and discouraging carbon-intensive ones. I assess how RGGI affected the size of each source in the electricity generation process over time. I find that RGGI sharply cut fossil fuel use, especially coal, by shifting generation to gas and nuclear rather than boosting renewables, underscoring the need for complementary policies to accelerate a truly renewable-driven energy transition.
@@
First, it contributes to the extensive DiD literature. Classical DiD focuses on scalar outcomes, estimating treatment effects on the mean, with surveys and applications summarized in Lechner (2011); Cafri et al. (2019); De Chaisemartin and d’Haultfoeuille (2023); Roth et al. (2023); Baker et al. (2025). Recent works have extended DiD to more complex settings: Callaway and Sant’Anna (2020) develops methods for staggered treatment adoption, Arkhangelsky et al. (2021) combines DiD with synthetic controls, and Manski and Pepper (2018); Rambachan and Roth (2020); Ban and Kédagni (2022) relax the parallel trends assumption. A second wave of research extends DiD to entire outcome distributions. Early contributions include quantile DiD Meyer et al. (1990) and the changes-in-changes (CiC) approach Athey and Imbens (2006), which identifies counterfactual distributions non-parametrically. Subsequent work has generalized these ideas to multivariate outcomes Torous et al. (2021), settings where identifying assumptions apply to cumulative distribution functions Havnes and Mogstad (2015); Roth and Sant’Anna (2021), copula-based approaches Callaway et al. (2018); Callaway and Li (2019); Ghanem et al. (2023), and methods using characteristic functions Bonhomme and Sauder (2011). However, these methods do not directly address categorical outcome distributions, which are probability mass functions. Graves et al. (2022) proposes a DiD for categorical outcomes while relying on parallel trends for proportions, which I explained earlier is not appropriate in general. Zhou et al. (2025) extend the difference-in-differences framework to non-Euclidean data, including probability mass functions, by employing Fréchet means and geodesic transport. While their approach is mathematically elegant, the identifying assumptions are primarily geometric and offer limited economic interpretability. In contrast, my framework delivers both counterfactual quantities and probability mass functions that retain economic intuition while incorporating a geometric perspective specifically designed for categorical outcomes. Another related paper is Tchetgen Tchetgen et al. (2024), who propose a general DiD framework applicable to count data in the binomial setting. My analysis differs by focusing on the multinomial case, allowing for richer categorical structures and more general forms of distributional change. Second, the paper contributes to the compositional data analysis (CoDA) literature. CoDA, pioneered by Aitchison (1982, 1990, 1992, 2002), provides tools for analyzing data constrained to the simplex. Subsequent developments Egozcue et al. (2003); Billheimer et al. (2001); Barceló-Vidal et al. (2001) refine transformations and models that respect the geometric structure of compositional data. More recently, Arnold et al. (2020) connected CoDA to causal inference. I also combine compositional data methods with econometric identification strategies to analyze causal effects when outcomes are probability mass functions, providing a bridge between DiD and compositional statistics.